### v2

# Matthias von Davier

Abstract
This article presents a corrected version of the Satterthwaite (1941, 1946) approximation for the degrees of freedom of a weighted sum of independent variance components. The original formula is known to yield biased estimates when component degrees of freedom are small. The correction, derived from exact moment matching, adjusts for the bias by incorporating a factor that accounts for the estimation of fourth moments. We show that Kish’s (1965) effective sample size formula emerges as a special case when all variance components are equal and component degrees of freedom are ignored. Simulation studies demonstrate that the corrected estimator closely matches the expected degrees of freedom even for small component sizes, while the original Satterthwaite estimator exhibits substantial downward bias. Additional applications are discussed, including jackknife variance estimation, multiple imputation total variance, and the Welch test for unequal variances.
1 Introduction
The estimation of effective degrees of freedom for complex variance estimators is a recurring problem in statistics, particularly when combining independent variance components with different degrees of freedom. Satterthwaite (1941, 1946) provided a widely used approximation that expresses the degrees of freedom of a weighted sum of variance components as a function of the component variances and their degrees of freedom. This approximation underpins many statistical procedures, including the Welch test for two-sample comparisons with unequal variances, and variance estimation in complex survey designs. However, the Satterthwaite formula relies on large-sample approximations and can be severely biased when the component degrees of freedom are small or the number of components is limited (Johnson & Rust, 1992).
In parallel, Kish (1965) introduced the concept of effective sample size for weighted samples, which adjusts the nominal sample size to account for unequal weighting. Kish’s formula, derived under the assumption of equal variances across observations, has become a standard tool in survey sampling. Although developed in different contexts, both Satterthwaite’s and Kish’s estimators share a common structure and can be linked through a simplified Satterthwaite expression when component variances are homogeneous.
This article has three main objectives. First, we derive a corrected Satterthwaite degrees-of-freedom estimator that mitigates the bias for small component degrees of freedom. The correction follows from a careful handling of the expected fourth moments of the component variance estimates, leading to an improved formula that reduces to the original Satterthwaite expression as component degrees of freedom increase. Second, we show that Kish’s effective sample size is a special case of the Satterthwaite estimator when all variance components are equal and the component degrees of freedom are effectively ignored. Third, we present simulation results comparing the original Satterthwaite estimator, the corrected version, and Kish’s estimator under various conditions, demonstrating the bias reduction achieved by the correction. Finally, we illustrate applications of the corrected formula in jackknife variance estimation, multiple imputation total variance, and the Welch test.
The article is organized as follows. Section 2 reviews Satterthwaite’s original derivation and presents the corrected formula. Section 3 discusses Kish’s effective sample size and its relationship to Satterthwaite’s estimator. Section 4 provides simulation evidence comparing the three estimators. Section 5 gives examples of the corrected estimator beyond effective sample size, including jackknifing, multiple imputation, and the Welch test. Section 6 concludes with a summary of recommendations.
2 Satterthwaite’s (1941, 1946) Approximate d.f.
Satterthwaite (1941, 1946) examines the approximate degrees of freedom for complex variance estimates. That is, we have weights $w_{1},\dots,w_{k}$and variance components $S_{1}^{2},\dots S_{k}^{2}$, each with degrees of freedom $\nu_{k}$, so that
$\frac{\nu_{k}S_{k}^{2}}{\sigma_{k}^{2}}\sim\chi_{\nu_{k}}^{2}$
that is,
$E\left(\frac{\nu_{k}S_{k}^{2}}{\sigma_{k}^{2}}\right)=\nu_{k}\>\mathrm{and}\>Var\left(\frac{\nu_{k}S_{k}^{2}}{\sigma_{k}^{2}}\right)=2\nu_{k}.$
The approximate d.f. are developed for a complex variance estimate
$S^{2}=\sum_{k=1}^{K}w_{k}S_{k}^{2}$
which is a linear combination of independent variance components. Satterthwaite (1941, 1946) argues that with independent $X_{k}$ and reasonably large d.f. for each of the variance components $S_{k}^{2}$, one may assume that
$S^{2}\sim\chi_{\nu^{*}}^{2}$
and that
$d.f.\left(S^{2}\right)=\nu^{*}\approx\frac{\left(\sum_{i}w_{i}S_{i}^{2}\right)^{2}}{\sum_{i}w_{i}^{2}\frac{S_{i}^{4}}{\nu_{k}}}.$
2.1 Proof of Satterthwaite’s Formula and its Improvement
The formula provided by Satterthwaite (1941, 1946) is widely used, but it is known that it produces biased estimated (e.g. Johnson & Rust, 1992) for small component degrees of freedom $\nu_{k}$. In this section, we derive the expected value the approximation is based on and provide a tighter approximation suitable for small $\nu_{k}$ and $K$, while it converges to the original for large $K$ and large $\nu_{k}$.
This section starts with reviewing some of the properties of $\chi^{2}$distributed variables that are needed to derive the main result.
2.1.1 Some Properties of $\chi^{2}$ Distributed Random Variables
Note that for any $\chi^{2}$ distributed variable with $\nu$ degrees of freedom we have
$\frac{\nu^{2}Var\left(S^{2}\right)}{\sigma^{4}}=2\nu\rightarrow\nu=2\frac{\left[\sigma^{2}\right]^{2}}{Var\left(S^{2}\right)}$
and
$\frac{\nu^{2}Var\left(S^{2}\right)}{\sigma^{4}}=2\nu\rightarrow Var\left(S^{2}\right)=2\frac{\left[\sigma^{2}\right]^{2}}{\nu}.$
Also, we have
$Var\left(S^{2}\right)+E\left(S^{2}\right)^{2}=E\left(S^{4}\right)=2\frac{\sigma^{4}}{\nu}+\sigma^{4}=\sigma^{4}\left(\frac{2}{\nu}+1\right)=\sigma^{4}\left(\frac{2+\nu}{\nu}\right)$
and hence
$\sigma^{4}=E\left(S^{4}\right)\frac{\nu_{S}}{\nu_{S}+2}.$
$Var\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)=\sum_{k=1}^{K}w_{k}^{2}Var\left(S_{k}^{2}\right)=\sum_{k=1}^{K}w_{k}^{2}E\left(S_{k}^{4}\right)-\sum_{k=1}^{K}w_{k}^{2}E^{2}\left(S_{k}^{2}\right)$
and further
$2\frac{\sigma^{4}}{\nu^{*}}=\sum_{k=1}^{K}w_{k}^{2}2\frac{\sigma_{k}^{4}}{\nu_{k}}.$
2.1.2 Derivation of the Satterthwaite Formula
Since the $S_{k}^{2}$ are independent we have $Var\left(S^{2}\right)=\sum_{k}w_{k}^{2}Var\left(S_{k}^{2}\right)$ and by using the equivalence $\nu=2\frac{\left[\sigma^{2}\right]^{2}}{Var\left(S^{2}\right)}\leftrightarrow Var\left(S^{2}\right)=2\frac{\left[\sigma^{2}\right]^{2}}{\nu}$ for $S^{2}$and $S_{k}^{2}$ we obtained
$2\frac{\sigma^{4}}{\nu^{*}}=\sum_{k=1}^{K}w_{k}^{2}2\frac{\sigma_{k}^{4}}{\nu_{k}}$
which, by solving for $\nu^{*},$provides
$\nu^{*}=\frac{\sigma^{4}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\sigma_{k}^{4}}{\nu_{k}}}$
Then, an estimate of $\nu^{*}$ is obtained by using the equivalence
$\sigma^{4}=E\left(S^{4}\right)\frac{\nu}{\nu+2}\approx S^{4}$
both for $\sigma_{k}^{4}=\left(\sigma_{k}^{2}\right)^{2}=\left[E\left(S_{k}^{2}\right)\right]^{2}$ and $\sigma^{4}.$ This assumes that the components degrees of freedom are sufficiently large so that $\frac{\nu}{\nu+2}\approx 1$ and that $S^{4}\approx E\left(S^{4}\right)$ which also is implied by $\nu\rightarrow\infty.$
Then, plugging in $S^{4}$ for $\sigma^{4}$, as well as $S_{k}^{2}$ for $\sigma_{k}^{2}$ yields
$\nu^{*}\approx\frac{\left[\sum_{k}w_{k}S_{k}^{2}\right]^{2}}{\sum_{k}w_{k}^{2}\frac{\left[S_{k}^{2}\right]^{2}}{\nu_{k}}}.$
Satterthwaite points out this approximation may not work well for small $K,\nu_{k}$.
2.1.3 Some Equivalencies
Here, it is important to note that
$v_{satt}^{*}=\frac{\left[\sum_{k}w_{k}\sigma_{k}^{2}\right]^{2}}{\sum_{k}w_{k}^{2}\frac{\left[\sigma_{k}^{2}\right]^{2}}{\nu_{k}}}=\frac{\left[\sum_{k}cw_{k}\sigma_{k}^{2}\right]^{2}}{\sum_{k}\frac{\left[cw_{k}\sigma_{k}^{2}\right]^{2}}{\nu_{k}}}$
for any $c>0.$ Choosing
$c=K\left[\sum_{k}w_{k}\sigma_{k}^{2}\right]^{-1}$
and setting $\overline{w\sigma^{2}}=\frac{1}{K}\sum_{k}w_{k}\sigma_{k}^{2}$ yields
$\nu_{satt}^{*}=\frac{\left[\sum\frac{w_{k}\sigma_{k}^{2}}{\overline{w\sigma^{2}}}\right]^{2}}{\sum_{k}\frac{1}{\nu_{k}}\left[\frac{w_{k}\sigma_{k}^{2}}{\overline{w\sigma^{2}}}\right]^{2}}=\frac{K^{2}}{\sum_{k}\frac{1}{\nu_{k}}\left[\frac{w_{k}\sigma_{k}^{2}}{\overline{w\sigma^{2}}}\right]^{2}}$
which is $K$ times the harmonic mean of the $q_{k}=\nu_{k}\left[\frac{\overline{w\sigma^{2}}}{w_{k}\sigma_{k}^{2}}\right]^{2}$
$\nu_{satt}^{*}=K\cdot H\left(q_{k}\right)=K\cdot H\left(\nu_{k}\left[\frac{\overline{w\sigma^{2}}}{w_{k}\sigma_{k}^{2}}\right]^{2}\right).$
This section shows that the neither the Satterthwaite estimator, nor its improved version, depend on the average value $\overline{w\sigma^{2}}$ of the weighted variances $w_{i}\sigma_{k}^{2}$.
2.1.4 Improvement for Small $\nu_{k}$ and Small $K$
While Satterthwaite assumes that $K,\nu_{k}>>1$ and hence, for all instances, his work implicitly uses $\frac{\nu}{\nu+2}\approx 1$, and then replaces $\sigma_{k}^{2}$ by $S_{k}^{2}$ directly, without the factor for estimation of $\nu_{satt}^{*}$, von Davier (2025a,b) points out that for small $K,\nu_{k}$ a correction is needed. Recall that we have shown above that
$\hat{\sigma}_{k}^{4}\approx S_{k}^{4}\frac{\nu_{k}}{\nu_{k}+2}\;\mathrm{and}\;\sigma^{4}\approx S^{4}\frac{\nu^{*}}{\nu^{*}+2}$
which provide the necessary correction. It follows that
$\nu^{*}\approx\frac{\frac{\nu^{*}}{\nu^{*}+2}\left[\sum_{k}w_{k}S_{k}^{2}\right]^{2}}{\sum_{k}w_{k}^{2}\frac{\left[S_{k}^{2}\right]^{2}}{\nu_{k}+2}}\leftrightarrow\nu^{*}\approx\frac{\left[\sum_{k}w_{k}S_{k}^{2}\right]^{2}}{\sum_{k}w_{k}^{2}\frac{\left[S_{k}^{2}\right]^{2}}{\nu_{k}+2}}-2.$
2.2 Some Results on the Improved Effective D.F.
The following tables show results for sums of $K$ simple variance estimates $S_{k}^{2}$ with
$\frac{\nu_{k}S_{k}^{2}}{\sigma_{k}^{2}}$
$\chi^{2}$ distributed. The ideal case where $\forall k,j:\sigma_{k}^{2}=\sigma_{j}^{2},\nu_{k}=\nu_{j},w_{k}=w_{j}$ was used to be able to compare the original Satterthwaite formula to the improved estimator, and against the expected value
$E\left(\nu^{*}\right)=K\overline{\nu}$
that follows in this case since each component $S_{k}^{2}$ has the same distribution.
Table 1: Simulation results for $K=2$ and $K=4$
$K$ dfdf $\bar{\nu}_{\text{unc}}$ SD unc $\bar{\nu}_{\text{corr}}$ SD corr $K\times\bar{\nu}$
2 1 1.410 0.346 2.229 1.039 2
2 2 3.135 0.640 4.270 1.281 4
2 4 6.837 1.044 8.256 1.566 8
2 8 14.577 1.494 16.222 1.868 16
2 16 30.348 1.938 32.141 2.181 32
2 32 62.199 2.282 64.086 2.425 64
4 1 2.198 0.637 4.595 1.912 4
4 2 5.320 1.202 8.640 2.403 8
4 4 12.400 1.974 16.600 2.961 16
4 8 27.567 2.852 32.459 3.565 32
4 16 58.908 3.567 64.272 4.013 64
4 32 122.482 4.113 128.138 4.370 128
Table 2: Simulation results for $K=8$ and $K=16$
$K$ dfdf $\bar{\nu}_{\text{satt}}$ SD unc $\bar{\nu}_{\text{corr}}$ SD corr $K\times\bar{\nu}$
8 1 3.656 1.023 8.969 3.070 8
8 2 9.473 2.014 16.946 4.029 16
8 4 23.171 3.296 32.757 4.944 32
8 8 53.198 4.720 64.498 5.900 64
8 16 115.997 5.736 128.497 6.453 128
8 32 243.096 6.552 256.290 6.962 256
16 1 6.443 1.608 17.330 4.825 16
16 2 17.586 3.180 33.172 6.360 32
16 4 44.642 5.226 64.963 7.838 64
16 8 104.373 7.278 128.466 9.097 128
16 16 229.741 8.776 256.458 9.873 256
16 32 484.036 9.730 512.289 10.338 512
Table 3: Simulation results for $K=32$ and $K=64$
$K$ dfdf $\bar{\nu}_{\text{satt}}$ SD unc $\bar{\nu}_{\text{corr}}$ SD corr $K\times\bar{n}$
32 1 11.938 2.503 33.815 7.509 32
32 2 33.825 4.924 65.650 9.847 64
32 4 87.394 7.914 129.091 11.870 128
32 8 207.089 10.860 256.861 13.575 256
32 16 457.054 13.031 512.186 14.659 512
32 32 965.813 14.191 1024.177 15.078 1024
64 1 22.725 3.757 66.175 11.270 64
64 2 65.904 7.259 129.807 14.517 128
64 4 172.991 11.588 257.486 17.381 256
64 8 412.037 15.757 513.047 19.696 512
64 16 912.294 18.557 1024.331 20.876 1024
64 32 1929.855 20.471 2048.471 21.750 2048
The obvious conclusion is that the correction of the Satterthwaite effective sample size is necessary even for large component degrees of freedom. The case where $K=64$ and $\nu_{k}=\overline{\nu}=32$ shows this clearly. The average $df$ for the corrected estimate is $\overline{\nu_{corr}^{*}}=2048.5\approx 2028=K\times\overline{\nu}>\overline{\nu_{satt}^{*}}=1929.9$. The other end of the table shows for $K=2$ and $df=1$ that $\overline{\nu_{corr}^{*}}=2.23\approx 2=K\times\overline{\nu}>1.41=\overline{\nu_{satt}^{*}}.$
3 Kish’s (1965) Effective Sample Size Estimate
To start with the similarities, both Kish’s (1965) formula for the effective sample size for an estimator of the variance of the mean and the Satterthwaite (1941) estimator of effective degrees of freedom of a complex variance estimate conceptual similarities.
[…]
$N_{R}=\sum_{i=1}^{N}w_{i}=\left\|\left\{1,\dots,N\right\}\backslash R\right\|$
which can be considered the effective sample size for a weighted sample where some cases have unit weights ($w_{i}=1$), and the rest has vanishing ($w_{r}=0$) weights.
In more general terms, the weights $w_{i}$ may be real valued and positive, so that their relative variance
$relvar\left(\boldsymbol{w}\right)=\frac{1}{K}\sum_{k=1}^{K}\left(\frac{w_{k}}{\overline{w}}-1\right)^{2}$
with $\overline{w}=\frac{1}{K}\sum_{k}w_{k}$ does not vanish.
Consider a random variable $Y$ with finite mean $\mu$ and finite variance $\sigma^{2}$. Let $y_{i}\in\mathbb{R},i=1,\dots,N$ denote the observed values in a sample of size $K$ and assume the sample requires weighting using the weights $w_{k}$. We are studying a weighted estimator of the sample mean using weights $w_{k}\in\mathbb{R}^{+}$, that is
$\overline{y}_{w}=M_{w}\left(y\right)=\frac{\sum_{k=1}^{K}w_{k}y_{k}}{\sum_{k=1}^{K}w_{k}}=\frac{1}{N}\sum_{k=1}^{K}\frac{w_{k}}{\overline{w}}y_{k}.$
Similarly, we can estimate the variance of $Y$
$Var_{w}\left(y\right)=\frac{\sum_{k=1}^{K}w_{k}\left(y_{k}-\overline{y}_{k}\right)^{2}}{N\overline{w}}=\frac{\sum_{k=1}^{K}w_{k}y_{k}^{2}}{N\overline{w}}-\left[\frac{\sum_{k=1}^{K}w_{k}y_{k}}{N\overline{w}}\right]^{2}=\overline{y_{w}^{2}}-\left(\overline{y}_{w}\right)^{2}.$
The weights $w_{k}$ can be chosen to reflect unequal sampling probability. For example, if the probability of sampling observation $i$ is given by $P_{i}>0,$the weights could be chosen as $w_{i}=P_{i}^{-1}$ to counter unequal probabilities of being included in the sample.
Another scenario is that the $y_{i}$ are statistics based on smaller samples taken from certain well-defined groups (sampling units, sub-populations, etc.). In that case, $y_{k}=f_{k}\left(y_{k1},\dots,y_{kN_{k}}\right)$ are (potentially also weighted) aggregates based on the within group observations $y_{k1},\dots,y_{kN_{k}}$. In this case, each of the groups may receive a different weight to account for and correct discrepancies between expected population proportion vs. sample sizes for each group.
Two things are worth noting here:
  1. 1.
    The variance estimate $Var_{w}\left(y\right)$ under non-uniform weights may not be the same as the variance calculated under simple random sampling $Var_{1}\left(y\right)$ (where every observation is ’worth the same’, as all $w_{i}=1$) due to the weights being used.
  2. 2.
    The quality of the variance estimator $Var_{w}\left(y\right)$, may differ from the quality of the unweighted estimate $Var_{1}\left(y\right).$ The term quality is chosen to be deliberately vague here. One well known issue with weighting is that it improve one quality (reduce bias) at the cost of another (increase uncertainty).
[…]
$Var\left(\overline{y}_{w}\right)=Var\left(\frac{\sum_{k=1}^{K}w_{k}y_{k}}{\sum_{k=1}^{K}w_{k}}\right)=\frac{\sum_{k=1}^{K}w_{k}^{2}Var\left(y_{k}\right)}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}},$
which follows when assuming that the $y_{k}$ are independent. When we additionally assume that $\forall k:Var\left(y_{k}\right)=\sigma^{2}$, we obtain
$Var\left(\overline{y}_{w}\right)=\frac{\sum_{k=1}^{K}w_{k}^{2}\sigma^{2}}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}}=\sigma^{2}\frac{\sum_{k=1}^{K}w_{k}^{2}}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}}=\frac{\sigma^{2}}{N}\frac{\overline{w^{2}}}{\overline{w}^{2}}.$
Using the same notation as above, Kish’s (1965) estimate of effective sample size is given by
$n_{eff}\approx\frac{\left(\sum_{k=1}^{K}w_{k}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}}\rightarrow Var\left(\overline{y}_{w}\right)=\frac{\sigma^{2}}{n_{eff}}.$
Across the literature, there is another expression that is often given for the effective sample size. This expression is based on the relvariance, which stands for relative variance, of the weights. It is the variance of the weights divided by the average of the weights. That is
$D_{eff}=\left[1+cv^{2}\left(\boldsymbol{w}\right)\right]$
where
$1+cv^{2}\left(\boldsymbol{w}\right)=1+Var\left(\frac{\boldsymbol{w}}{\overline{w}}\right)=1+\frac{1}{N}\sum_{i=1}^{N}\left(\frac{w_{i}}{\overline{w}}-1\right)^{2}$
since $M\left(\boldsymbol{w}\right)=\overline{w}$. Then we have
$1+cv^{2}\left(\boldsymbol{w}\right)=\frac{1}{N\overline{w}^{2}}\sum_{k=1}^{K}w_{k}^{2}=N\frac{\sum_{k=1}^{K}w_{k}^{2}}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}}=\frac{N}{n_{eff}}=D_{eff}.$
4 Satterthwaite’s Formula implies Kish’s $n_{eff}$
For a sample $X_{1},...,X_{K}$ with weights $w_{1},\dots,w_{K}$ of $K$ observations, consider
$\nu^{*}\approx\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{S_{k}^{4}}{\nu_{k}}}$
which is the well known Satterthwaite estimator of the d.f. for the variance estimate
$S^{2}=\sum_{k=1}^{K}w_{k}S_{k}^{2}.$
The Kish (1965) approximation of the effective sample size can obtained by assuming that all variance components are identical, i.e., $\forall i,j:S_{i}^{2}=S_{j}^{2}=S_{0}^{2}$. In that case, we obtain
$\nu^{*}\approx\frac{\left(\sum_{k=1}^{K}w_{k}S_{0}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\left[S_{0}^{2}\right]^{2}}=\frac{\left[S_{0}^{2}\right]^{2}\left(\sum_{k=1}^{K}w_{k}\right)^{2}}{\left[S_{0}^{2}\right]^{2}\sum_{k=1}^{K}w_{k}^{2}}=n_{eff}.$
In this case, no small $K,\nu_{k}$ is needed as the assumption made by Kish eliminates the need to insert estimates of $E\left(S^{4}\right).$The result turns out to be a simplified Satterthwaite-type estimator of degrees of freedom if all component variances are known to be the same, and only the weights are unequal.
4.1 Comparing Kish, Satterthwaite, and the Corrected Version
[…]
Table 4: Results for different weights (random $w_{k}\sim N(1,0.3)$). Columns: mean Kish, mean Satterthwaite (uncorrected), mean corrected Satterthwaite, total sample size $K\overline{\nu}$, and ratios relative to $K$ or $K\overline{\nu}$.
[…]
16 1 14.74 6.16 16.49 16 0.92 0.39 1.03
16 5 14.74 55.36 75.50 80 0.92 0.69 0.94
16 50 14.75 712.31 738.81 800 0.92 0.89 0.92
16 500 14.75 7351.23 7378.63 8000 0.92 0.92 0.92
32 1 29.43 11.28 31.83 32 0.92 0.35 0.99
32 5 29.43 107.84 148.98 160 0.92 0.67 0.93
32 50 29.42 1417.19 1471.87 1600 0.92 0.89 0.92
32 500 29.43 14659.25 14715.88 16000 0.92 0.92 0.92
64 1 58.78 21.27 61.81 64 0.92 0.33 0.97
64 5 58.78 212.92 296.08 320 0.92 0.67 0.93
64 50 58.81 2830.18 2941.39 3200 0.92 0.88 0.92
64 500 58.79 29280.13 29395.25 32000 0.92 0.92 0.92
Table 5: Results for equal weights (all $w_{k}=1$). Columns: mean Kish, mean Satterthwaite (uncorrected), mean corrected Satterthwaite, total sample size $K\overline{\nu}$, and ratios relative to $K$ or $K\overline{\nu}$.
[…]
16 1 16.00 6.45 17.36 16 1.00 0.40 1.08
16 5 16.00 59.31 81.03 80 1.00 0.74 1.01
16 50 16.00 771.06 799.90 800 1.00 0.96 1.00
16 500 16.00 7970.16 8000.04 8000 1.00 1.00 1.00
32 1 32.00 11.93 33.78 32 1.00 0.37 1.06
32 5 32.00 116.40 160.96 160 1.00 0.73 1.01
32 50 32.00 1540.29 1599.91 1600 1.00 0.96 1.00
32 500 32.00 15938.14 15999.89 16000 1.00 1.00 1.00
64 1 64.00 22.75 66.24 64 1.00 0.36 1.04
64 5 64.00 230.86 321.20 320 1.00 0.72 1.00
64 50 64.00 3078.86 3200.01 3200 1.00 0.96 1.00
64 500 64.00 31874.41 31999.90 32000 1.00 1.00 1.00
It can be seen that with increasing $\overline{\nu}\rightarrow\infty$ the Satterthwaite and the Kish estimates converge. This is expected as Kish’s estimator is obtained by ignoring the component-wise degrees of freedom $\nu_{k}$. Therefore, Kish’s approximation should only be applied when it is known that all components estimate the same variance, and in addition, it is known that the effect of estimating the component variances can be ignored.
5 Examples Beyond Effective Sample Size
5.1 Jackknifing and Balanced Repeated Replications (BRR)
In the case of single component d.f. $\nu_{k}=1$ and large(-ish) $K$ representing the (non-trivial) jackknife zones we have no weights to account for and obtain
$\nu^{*}\approx\frac{3\left(\sum_{k=1}^{K}\left[T_{\cdot}-T_{k}\right]^{2}\right)^{2}}{\left[T_{\cdot}-T_{k}\right]^{4}}-2$
since $\nu_{k}+2=3$ for all $k.$ Here, the $T_{k}$ are the pseudo-values calculated under the jackknife scheme of dropping units, and $T_{\cdot}=\sum_{k}T_{k}$. In the case of the BRR the $T_{k}$ are the balanced replicated estimates (of the half samples) obtained by adjusting the weights of the pairs and recalculating the statistics for each zone $K$.
5.2 Total Variance under Multiple Imputations
The total variance of an estimator when imputations is a prime example of a weighted variance estimate. The total variance as defined by Rubin (1987) is given by
$Var\left(total\right)=Var\left(sampling\right)+\frac{M+1}{M}Var\left(imputation\right)$
where the sampling variance may be estimated according to some resampling scheme (Johnson & Rust, 1992) and the imputation variance is the sample standard deviation across $M$ imputation based calculations of the same statistic.
In this case, the following weights are used:
$\left(w_{1},w_{2}\right)=\left(1,\frac{M+1}{M}\right)$
and the corrected equation can be directly applied.
5.3 A Corrrected Welch Test $d.f.$
The modified Welch (1947) test uses the d.f. $\nu_{1}=N_{1}-1$ and $\nu_{2}=N_{2}-1$ of each of the variances of two samples to pool the variance. The Welch-Satterthwaite formula used in this case can be adjusted using the corrected Satterthwaite formula derived above. Then we have
$\nu^{*}\approx\frac{\left(\frac{1}{N_{1}}S_{1}^{2}+\frac{1}{N_{2}}S_{2}^{2}\right)^{2}}{\frac{S_{1}^{4}}{N_{1}^{2}\left(v_{1}+2\right)}+\frac{S_{1}^{4}}{N_{2}^{2}\left(v_{2}+2\right)}}-2$
where $S_{k}^{2}$ are the sample variances and weights
$w_{k}=\frac{1}{N_{k}}$
in this case, for the d.f. $\nu_{*}$ of the sample size weighted pooled variance used in the Welch test.
6 Conclusion
The approaches presented here were developed under different constraints. The Kish (1965, 1992) effective sample size puts the strongest constraints on the estimator, as it has been developed as tool for a simple approximation when only the weights are known and all components of a variance estimate can be assumed to have the same expectation, and each component is a well estimated variance with moderate to large degrees of freedom.
The Satterthwaite (1941, 1946) estimate of the effective degrees of freedom was developed for weighted sums of variance components with varying degrees of freedom. This is a much more general approach than the simple approxiamtion provided by Kish (1965, 1992), but lacks the ability to generate useful estimates when the component degrees of freedom, $\nu_{k},$are small.
The corrected Satterthwaite effective degrees of freedom is based on von Davier’s (2025a/b) research and produces estimates that are much closer to the expected value, when simulating data using the the ideal case, $K\overline{\nu}$, also for small component $\nu_{k}$. It is calculated as
$\nu^{*}\approx\frac{\left[\sum_{k}w_{k}S_{k}^{2}\right]^{2}}{\sum_{k}w_{k}^{2}\frac{\left[S_{k}^{2}\right]^{2}}{\nu_{k}+2}}-2$
and with growing component degrees of freedom $\nu_{k}$ and components $K$ it approaches the original Satterthwaite expression, while for small $\nu_{k}$ it avoids the severe bias found with the original formula.